OpenStudy (anonymous):

Find y in \[y=\int\limits_{0}^{\infty}\frac{1}{x^5(e^{\frac{k}{x}}-1)}dx\]

OpenStudy (anonymous):

integrate by parts?

OpenStudy (anonymous):

trig subs might work, but x would need to be like ln(tan^2x)^-1 making it overly complicated once you need to solve back for the original answer, and finding the derivative not sure

OpenStudy (anonymous):

\[\frac{d(e^{k/x}-1)^{-1}}{dx}=\frac{-1}{(e^{k/x}-1)^{2}}(ke^{k/x})(\frac{-1}{x^2})\] If you use the exponetial in parts as the bit that you differentiate, it will generate new problems. If you use \[1/x^5\], the 5 will grow.

OpenStudy (anonymous):

So parts looks hopeless.

OpenStudy (anonymous):